Using the equivalence point volumes determined from the expanded titration graph or the first derivative graph, the volume of acetic acid titrated (pipet volume), and the concentration of your diluted NaOH standard solution, calculate the concentration of the unknown acid solution.

Equivalence point volume= 18.5

Concentration of diluted NaOH= 0.19709

Volume acetic acid= 5 mL (diluted in 70 mL dH20 but I guess it does not matter?)
Answer #1

concentration of unkown acid can be calculated via:

mmol of acid = mmol of base

so

mmol of base = MV = 0.19709*18.5 = 3.646165 mmol of OH-

so

mmol of H+ = 3.646165 mmol

therefore

initial volume of acid:

V = 5 mL

[acid] = mmol/mL = 3.646165/5 = 0.729233 M of acid